### v1

Abstract
Fast and accurate simulation of soft tissue deformation is a critical factor for surgical robotics and medical training. In this paper, we introduce a novel physics-informed neural simulator that approximates soft tissue deformations in a realistic and real-time manner. Our framework integrates Kelvinlet-based priors into neural simulators, making it the first approach to leverage Kelvinlets for residual learning and regularization in data-driven soft tissue modeling. By incorporating large-scale Finite Element Method (FEM) simulations of both linear and nonlinear soft tissue responses, our method improves neural network predictions across diverse architectures, enhancing accuracy and physical consistency while maintaining low latency for real-time performance. We demonstrate the effectiveness of our approach by performing accurate surgical maneuvers that simulate the use of standard laparoscopic tissue grasping tools with high fidelity. These results establish Kelvinlet-augmented learning as a powerful and efficient strategy for real-time, physics-aware soft tissue simulation in surgical applications.
Keywords:
Neural Simulators Soft Tissue Deformation Kelvinlet
1 Introduction
Soft tissue deformation modeling is a critical need in computational biomechanics for surgical simulation, image registration, and robotic-assisted interventions. Accurate tissue simulation supports surgical planning, image-guided predictions of soft-tissue changes, and precise haptic feedback for robotics. Traditional soft tissue deformation modeling relies on physics-based methods like Finite Element (FE) models, which offer high accuracy but are computationally expensive for real-time use [27]. Alternative models, including mass-spring systems [17], mesh-free methods [2], and particle-based methods [14], approximate tissue behavior, but require substantial computation and struggle to achieve realistic modeling.
Recent advances in deep learning have enabled efficient, data-driven approximations of soft tissue deformations. CNNs, GNNs, and Transformers [20] can learn complex deformation patterns from simulations or real-world data, enabling real-time tissue simulation. However, they face two key limitations: 1) poor generalization to unseen force distributions and boundary conditions, leading to unreliable extrapolation, and 2) the risk of producing physically implausible deformations, particularly when trained on limited datasets.
To overcome these limitations, physics-informed machine learning integrates fundamental physical laws into neural network training, offering a promising approach to improving accuracy and generalization. Physics-Informed Neural Networks (PINNs) [22] enforce differential equations via soft constraints, while hybrid methods like differentiable FEM solvers [25] and coarse-to-fine neural surrogates [12] aim to balance accuracy and efficiency. However, these approaches are computationally demanding, particularly for large soft-tissue meshes, and require careful consideration of numerical stability and differentiability.
We propose a physics-informed neural network framework that leverages Kelvinlets [6] to improve the accuracy and physical consistency of soft tissue deformation modeling. As a closed-form Green’s function solution to the Navier-Cauchy equations, Kelvinlets provide an efficient and analytically tractable prior. In this paper, we introduce two hybrid approaches: 1) using Kelvinlets as a baseline and learning the residual (modeling error) via a neural network, and 2) incorporating Kelvinlets as a regularizer for a neural network predicting the full deformation field. We generated a finite element dataset of liver deformations, using a linear model without preload and a nonlinear model with gravity-induced preload. We demonstrate that our proposed Neural-Augmented Kelvinlet outperforms purely data-driven models in generalization, physical realism, and learning efficiency, enabling real-time realistic modeling of soft-tissue deformations with multiple graspers.
Figure 1: Visualization of our Neural Augmented Simulator with realistic textures.
Contributions. Our work makes the following key contributions [^1] Repository available at https://github.com/Anon92373/Neural-Kelvinlet.:
  1. 1.
    We present a dataset of 20,800 FEM solutions for data-driven soft tissue modeling, with 10,400 linear and 10,400 nonlinear simulations. It includes 5,000 single-grasper and 5,400 multi-grasper cases.
  2. 2.
    We provide a GPU-accelerated Kelvinlet PyTorch implementation that enables efficient training and real-time inference, making physics-informed deformation modeling accessible for interactive applications.
  3. 3.
    We introduce two hybrid approaches that integrate Kelvinlets into data-driven soft tissue deformation modeling and conduct extensive experiments to demonstrate their superiority over purely physics-informed and data-driven baselines across diverse neural architectures.
2 Kelvinlets for Soft Tissue Deformation
2.1 Navier-Cauchy Equations and Green’s Function Solution
The displacement field $\mathbf{u}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ of an isotropic, homogeneous elastic medium under an applied body force density $\mathbf{f}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ satisfies the Navier-Cauchy equations:
$\mu\nabla^{2}\mathbf{u}+\frac{\mu}{1-2\nu}\nabla(\nabla\cdot\mathbf{u})+\mathbf{f}=0,$ (1)
where $\mu>0$ is the shear modulus, and $\nu\in(0,0.5)$ is the Poisson’s ratio, ensuring a physically consistent material response. For a concentrated force $\mathbf{f}\delta(\mathbf{x}-\mathbf{x}_{s})$ applied at a source point $\mathbf{x}_{s}\in\mathbb{R}^{3}$, the Green’s function solution, known as the Kelvinlet, provides an analytical displacement field $\mathbf{u}:\mathbb{R}^{3}\setminus\{\mathbf{x}_{s}\}\to\mathbb{R}^{3}$ following:
$\mathbf{u}(\mathbf{x})=\frac{1}{4\pi\mu}\left(\frac{(1-\nu)I}{r}+\nu\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r^{3}}\right)\mathbf{f},$ (2)
where $r=\|\mathbf{x}-\mathbf{x}_{s}\|$ denotes the Euclidean distance from the source.
The classical Kelvinlet in Eq. (2) exhibits singularities at $\mathbf{x}_{s}$. To prevent numerical instabilities, we substitute $r$ with $r_{\varepsilon}=\sqrt{\|\mathbf{x}-\mathbf{x}_{s}\|^{2}+\varepsilon^{2}}$ for $\epsilon>0$, where the spatial extent of force distribution is controlled by $\varepsilon$, ensuring the solution is well-conditioned. Finally, a re-scaled solution with respect to the local displacement boundary condition $\mathbf{u}_{s}=\mathbf{u}(\mathbf{x}_{s})$ removes the dependence on both the force $\mathbf{f}$ and the modulus $\mu$, yielding the regularized Kelvinlet formulation used in this paper,
$\mathbf{u}_{\varepsilon}(\mathbf{x})=\frac{\varepsilon}{(5-6\nu)}\left(\frac{(3-4\nu)I}{r_{\varepsilon}}+\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r_{\varepsilon}^{3}}\right)\mathbf{u}_{s}.$ (3)
For multiple interaction points, the Kelvinlet model represents the total organ deformation as $\mathbf{u}(\mathbf{x})=\sum_{i}k_{i}\mathbf{u}_{\varepsilon,i}(\mathbf{x}),$ where the coefficients $k_{i}$ are optimized to satisfy displacement boundary conditions $\mathbf{u}(\mathbf{x}_{i})=\mathbf{u}_{i}$, via:
$\operatorname*{argmin}_{\mathbf{k}}(\sum_{i}\|\sum_{j}k_{j}\mathbf{u}_{\varepsilon,j}(\mathbf{x}_{i})-\mathbf{u}_{i}\|^{2})+\lambda\sum_{i}k_{i}^{2}.$ (4)
This ensures that the contributions of individual Kelvinlet functions collectively satisfy the prescribed displacements. The regularization $\lambda=0.001$ ensures a minimum energy solution. The total Kelvinlet deformation model, including the optimization, can be performed in less than 3 ms for two simultaneously engaged graspers using native GPU acceleration in pytorch on a mesh with $\sim 10\text{k}$ nodes.
2.2 Method 1: Kelvinlet-Based Residual Learning
[…]
$\mathcal{L}_{\text{residual}}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\|\mathbf{u}_{\text{true}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})-\tilde{\mathbf{u}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})\|^{2}_{L_{2}(\Omega)}\right],$ (5)
where $p$ represents the distribution of all possible pointwise interactions with the tissue, and $\Omega$ denotes the tissue domain. By focusing on learning only the correction beyond the Kelvinlet solution, this approach enhances efficiency and interpretability while leveraging a known analytical model as a strong prior.
2.3 Method 2: Kelvinlets as Regularization
Alternatively, a network can be trained to directly regress $\mathbf{u}_{\text{true}}$, producing an estimate $\hat{\mathbf{u}}$. To enforce physically consistent predicted deformations by the network, we leverage Kelvinlets as a regularization prior and write:
$\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right]+\lambda_{reg}\mathbb{E}_{(\mathbf{x}^{\prime}_{s},\mathbf{u}^{\prime}_{s})\sim q}\left[\left\|\mathbf{u}_{\epsilon}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right],$ (6)
where $\lambda_{\text{reg}}>0$ is the regularization coefficient. For clarity and brevity, we omit the dependencies $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})$ and $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})$ in the notation. Note that we explicitly separate $(\mathbf{x}_{s},\mathbf{u}_{s})$ from $(\mathbf{x}_{s}^{\prime},\mathbf{u}^{\prime}_{s})$ because, in practice, the two distributions $p$ and $q$ could differ, and also while Monte Carlo integration is used for both terms, the first term requires FEM simulations to obtain $\mathbf{u}_{\text{true}}$ (which is computationally expensive), whereas the second term only involves Kelvinlet evaluations (which are much faster). We will describe the distributions $p$ and $q$ used in our experiments in the following section.
For both proposed methods, when multiple graspers interact with the tissue, we employ the multi-point interaction solution of Kelvinlet and provide all interaction points along with their corresponding displacements to the network.
3 Large-Scale FEM Data
Our neural network framework trains on a large FE simulation dataset, focused on lifting liver lobes to expose the gallbladder during laparoscopic cholecystectomy. We construct an FE model from an abdominal CT image and its segmentation masks, converting them into a 3D linear tetrahedral mesh with 10,400 nodes and 44,045 tetrahedra using ‘gmsh’ [9].
Interactions of the liver and gallbladder with surrounding organs are modeled by prescribing boundary conditions. Those boundary conditions restrict regions of the liver’s exposed surface that are considered in the far field from the interaction location, as well as the distal end of the cystic ducts, arteries, and veins. FE simulations are solved using the open source getFEM package [24].
[…]
$\boldsymbol{E}=\frac{1}{2}\left(\nabla\mathbf{u}+\nabla\mathbf{u}^{T}+\nabla\mathbf{u}^{T}\nabla\mathbf{u}\right).$ (7)
In addition, this equation relates to the Cauchy-Green deformation tensor $\boldsymbol{C}$ by: $\boldsymbol{C}=2\boldsymbol{E}-\mathbf{I}$, where $\mathbf{I}$ is the identity tensor.
For our nonlinear simulations, we use a two-parameter Mooney-Rivlin model, where the strain energy density can be written as functions of the Cauchy-Green deformation tensor invariants ($\boldsymbol{I_{i}(C)}$, with $i=1,2,3$) [10]:
$\boldsymbol{W}=C_{01}\left(\boldsymbol{I_{1}(C)}\boldsymbol{I_{3}(C)}^{\frac{-1}{3}}-3\right)+C_{10}\left(\boldsymbol{I_{1}(C)}\boldsymbol{I_{3}(C)}^{\frac{-2}{3}}-3\right),$ (8)
where $C_{10}=1.62$ kPa and $C_{01}=1.97$ kPa are constitutive constants derived from ex-vivo porcine liver experiments [3].
For our linear simulations, we simulate an elastic response in the absence of gravitational load ($\boldsymbol{g}=0$). These solutions serve as baseline data for tissue deformation under small elastic input displacements, where the term $\nabla\mathbf{u}^{T}\nabla\mathbf{u}$ in Eq. (7) is neglected. Assuming isotropic material properties, the strain energy density $\boldsymbol{W}$ in terms of shear modulus $\mu$ and Poisson’s ratio $\nu$ expressed as:
$\boldsymbol{W}=\left(\frac{\mu\nu}{1-2\nu}\right)\mathrm{Tr}(\boldsymbol{E})^{2}+\mu\mathrm{Tr}(\boldsymbol{E}^{2}).$ (9)
In these linear simulations, we assign a Poisson’s ratio of 0.45 for all organs [4]. The shear modulus values are also set as follows: 0.03 kPa for the fascia [1], 0.34 kPa for the gallbladder [18], 0.69 kPa for the gallbladder wall and liver [11], and 1.07 kPa for the remaining organs [8].
3.0.1 Training Distribution, $p$.
Given a mesh partitioned into regions $\Omega_{i}$ informed by cholecystectomy surgery, nodes $\mathbf{x}_{s}$ are sampled from a mixture of Generalized von Mises-Fisher (vMF) distributions (i.e., radial densities on the surface mesh) centered in these regions:
$p(\mathbf{x}_{s})=\sum_{i}g(\Omega_{i})f_{\text{vMF}}(\mathbf{x}_{s}\mid\boldsymbol{\mu}_{i},\kappa_{i}),\vspace{-.1in}$ (10)
where $g(\Omega_{i})$ is the prior-informed probability of selecting $\Omega_{i}$, $\boldsymbol{\mu}_{i}$ indicates the center of $\Omega_{i}$, and $\kappa_{i}$ is the spread of the vMF, which is a function of the size of $\Omega_{i}$. The signals $u_{s}$ are independently sampled from a uniform distribution in a cuboid $\mathcal{C}$ defined by: $x\in(x_{1},x_{2}),y\in(y_{1},y_{2}),z\in(z_{1},z_{2})$, with volume: $V=(x_{2}-x_{1})(y_{2}-y_{1})(z_{2}-z_{1}).$ This approach ensures that node sampling is spatially structured according to surgical priors while allowing displacements to be isotropic in a bounded range.
For single grasping, ten regions $\Omega_{i}$ around the liver edge were selected, each with 250 samples. For multiple graspers, six pairwise combinations of five regions $\Omega_{i}$ were considered, with 30 samples per region. The displacement ranges are $x\in(x_{1},x_{2})=(-90,50)$ mm, $y\in(y_{1},y_{2})=(-90,50)$ mm, and $z\in(z_{1},z_{2})=(-60,90)$ mm, corresponding to the cube domain.
3.0.2 Distribution for Multiple Graspers.
For multiple interactions (e.g., two graspers), we define the joint distribution as $p(\mathbf{x}_{s}^{1},\mathbf{x}_{s}^{2})=p(\mathbf{x}_{s}^{1})p(\mathbf{x}_{s}^{2}\mid\mathbf{x}_{s}^{1})$, where $p(\mathbf{x}_{s}^{1})$ follows Eq. (10). To ensure that $\mathbf{x}_{s}^{2}$ is not sampled from the same region $\Omega_{k}$ as $\mathbf{x}_{s}^{1}$, we modify the region probabilities $g(\Omega_{i})$ by setting $g(\Omega_{k})=0$ and adjusting $g(\Omega_{i\neq k})$ based on surgical priors, and then sampling $\mathbf{x}_{s}^{2}$ from $p(\mathbf{x}_{s}^{2}\mid\mathbf{x}_{s}^{1})$.
3.0.3 Regularization Distribution, $q$.
For Kelvinlet regularization, interaction points are sampled uniformly from the surface nodes. Initial displacements are then drawn uniformly from a cone centered around the surface normal at $\mathbf{x}_{s}$, with a maximum opening angle of $\alpha_{\max}$.
4 Texture Rendering for Realistic Simulations
Our visualization pipeline combines real-time tissue deformation and high-fidelity rendering to create a realistic simulation (see Figure 1 and the supplementary videos). In a preprocessing stage, surface meshes are extracted from the segmented high-resolution volumetric mesh, assigned texture coordinates, and enriched with procedurally generated textures using Worley noise [26]. Interpolation weights are computed to map low-resolution deformations to the high-resolution mesh [16], optimized using a Bounding Volume Hierarchy (BVH) [5], and UV coordinates are transferred to the volumetric mesh. At runtime, the system dynamically updates surface meshes to accommodate added or removed tetrahedral elements, interpolates high-resolution vertex positions based on low-resolution deformations, smooths exposed internal structures, and recalculates surface normals. Rendering employs pre-integrated diffuse shading [19], triplanar texture projection for inner tissue surfaces [7], and post-processing effects within Unity to achieve realistic tissue appearance. The pipeline achieves near-real-time performance while preserving visual fidelity during dynamic deformations.
5 Experiments
5.0.1 Data.
Our dataset consists of large-scale FEM simulations of soft tissue deformation. Each node in the mesh has a feature vector: $\mathbf{f}=\left[\mathbf{x};\mathbf{u}_{s};a\right]\in\mathbb{R}^{7}$. Where $\mathbf{x}\in\mathbb{R}^{3}$ is the undeformed position, $\mathbf{u}_{s}\in\mathbb{R}^{3}$ is the initial displacement, and $a\in\{0,1\}$ indicates whether the node is actively influenced by external forces. This distinction helps differentiate constrained nodes from passively deforming ones. We create a dataset of 20,800 FEM solutions with 10,400 linear elastic simulations and 10,400 nonlinear material responses. Of these, 5,000 correspond to single manipulations and 5,400 to multiple grasper manipulations.
5.0.2 Neural Architectures.
We evaluate Kelvinlet-based priors using four neural architectures: U-Net [15], SAB [13] (set attention), GraphGPS [23] (a Graph Neural Network and Transformer hybrid), and PointNet [21] (a universal function approximator for sets). We modify PointNet and GraphGPS by reducing layers, U-Net by replacing convolutions with shared MLPs to form a DeepSet, and SAB by using two layers with an MLP regressor head.
Each model is trained separately on two datasets: one with individual graspers and another with multiple graspers interacting. For each dataset, models are trained under three settings: (1) a baseline without physics priors (i.e., evaluating solely the network), (2) residual learning predicting deviations from Kelvinlets, and (3) physics-informed regularization enforcing Kelvinlet constraints.
5.0.3 Hyperparameters.
We set shear modulus in Kelvinlet to $\mu=0.72$ kPa, Poisson’s ratio $\nu=0.45$, and regularization parameter $\epsilon=0.05$, to reflect soft tissue properties and ensure numerical stability. For linear and nonlinear simulations we use $\lambda_{\text{reg}}=1$ and $\lambda_{\text{reg}}=0.1$, respectively, to emphasize linearity of Kelvinlet.
5.1 Results and Analysis
Figure 2: Qualitative comparison of deformation predictions for the best (PointNet) and worst (GraphGPS) models in linear (top) and nonlinear (bottom) regimes.
To quantitatively assess model performance, we introduce the Deformation Capture Mean (DCM) metric, which measures how well a predicted deformation field approximates the ground-truth displacement field. The DCM is computed as:
$\text{DCM}=100\times\left(1-\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\frac{\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|_{L_{2}(\Omega)}}{\|\mathbf{u}_{\text{true}}\|_{L_{2}(\Omega)}}\right]\right).$ (11)
Tables 1 and 2 report DCMs across models and settings. Each experiment runs with three random seeds, reporting mean and standard deviation. Models perform better on linear cases, as nonlinear deformations introduce gravity effects, non-proportional stress-strain relationships, and greater sensitivity to perturbations. Notably, Kelvinlet priors consistently improve DCMs, despite being derived for linear elasticity, showing their effectiveness even in nonlinear regimes.
Table 1: DCM of models with/without Kelvinlet priors in linear simulations.
Model Indiv. Graspers Comb. Graspers
Base +Kelv. Res. +Kelv. Reg. Base +Kelv. Res. +Kelv. Reg.
U-Net[15] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11) 88.23 (0.08) 89.67 (0.09) 90.01 (0.08)
GraphGPS[23] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23) 87.33 (0.06) 87.92 (0.02) 88.01 (0.04)
PointNet[21] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09) 89.79 (0.07) 90.11 (0.07) 91.22 (0.07)
SAB[13] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04) 88.05 (0.09) 89.91 (011) 91.02 (0.04)
Kelvinlet 65.24 (0) 64.33 (0)
Table 2: DCM of models with/without Kelvinlet priors in nonlinear simulations.
Model Indiv. Graspers Comb. Graspers
Base +Kelv. Res. +Kelv. Reg. Base +Kelv. Res. +Kelv. Reg.
U-Net[15] 80.56 (0.10) 84.22 (0.09) 84.88 (0.11) 76.91 (0.05) 79.83 (0.08) 82.10 (0.21)
GraphGPS[23] 77.52 (0.17) 80.63 (0.13) 82.11 (0.09) 73.72 (0.15) 77.92 (0.10) 80.66 (0.16)
PointNet[21] 81.25 (0.22) 84.65 (0.19) 85.05 (0.15) 79.56 (0.18) 82.66 (0.16) 83.33 (0.11)
SAB[13] 80.75 (0.19) 84.44 (0.07) 84.75 (0.10) 78.88 (0.11) 81.50 (0.15) 82.83 (0.16)
Kelvinlet 58.88 (0) 51.67 (0)
Figure 3 illustrates the trade-off between inference time and accuracy for different models across linear and nonlinear simulations. The results indicate that Kelvinlet regularization consistently improves accuracy compared to the base models, as evidenced by the upward shift in solid markers relative to their transparent counterparts. However, this enhancement in accuracy does not increase computational cost, maintaining inference times within a comparable range. Additionally, models trained on multi-grasper data (black-edged markers) exhibit similar trends to their single-grasper counterparts, suggesting that Kelvinlet regularization generalizes well across different grasping conditions. All computations were performed on a single NVIDIA A6000 GPU, ensuring a consistent hardware environment for runtime measurements.
Figure 3: Log-scaled inference time vs. accuracy (DCM%) for various models in linear and nonlinear simulations. Transparent markers denote base models, while solid markers indicate Kelvinlet-regularized versions. Black-edged markers represent multi-grasper cases, while edge-free markers denote single-grasper cases. FEM and Kelvinlet baselines are included for reference.
6 Conclusion
We introduce Neural-Augmented Kelvinlets, a physics-informed approach for real-time soft tissue deformation modeling that enhances neural network predictions in accuracy, physical consistency, and generalization while maintaining real-time performance. We utilize Kelvinlets in two settings: as a residual learning framework to refine predictions beyond an analytical baseline and as a regularization term to enforce physically plausible deformations. Large-scale FEM simulations provide a benchmark for evaluating data-driven deformation models, demonstrating that Kelvinlet integration significantly improves performance, particularly in complex multi-grasper scenarios. Our GPU-accelerated PyTorch implementation enables real-time inference, making it well-suited for surgical robotics and medical training, while a high-fidelity rendering pipeline ensures realistic soft tissue simulations in real time.
Acknowledgment
This work was partially supported by the NSF CAREER Award #2339898 and the Wellcome Leap ‘Surgery: Assess/Validate/Expand (SAVE)’ program.